OpenStudy (anonymous):

Rewrite f(x) = x2 + 9x – 16 in vertex form. Identify the axis of symmetry, vertex, and whether the vertex is a maximum or minimum help please

OpenStudy (anonymous):

first coordinate of the vertex is always \[-\frac{b}{2a}\]which is also the axis of symmetry in your case it is \[-\frac{9}{2\times 1}=-\frac{9}{2}\]

OpenStudy (anonymous):

therefore to write it in "vertex form" you can write it as \[x^2+9x-16=(x+\frac{9}{2})^2+k\] to find \(k\) replace \(x\) in your original expression by \(\frac{9}{2}\)